OpenStudy (anonymous):

Grade Counts On This ; Help Please <3

OpenStudy (anonymous):

I. Interest is ____?. (1pt) a. The amount of money your bank pays you for the use of only your money. b. The amount you might pay for a loan. * c. A percentage rate. d. All of the above. II. When calculating interest , you need too ______? (1pt) a. Change the percentage to a decimal then multiply the decimal against the amount. b. Just multiply the percentage against the amount without changing it. c. Add the percentage to the amount. d. Subtract the percentage from the amount.* III. If you place $1,000.00 in a savings account with an interest rate of 3.25 % / Month , How much should you earn in interest at the end of the month?. a. $ 3.25 b. $ 32.50* c.$ 325.00 d.$ 0.33 VIII. If you placed $ 1,000.00 in a savings account with an interest rate of 4.5 % versus an account with 3.75 % interest , how much more could you earn in interest for that month? a. $1.25 more b. $12.50 more * c. $7.50 more d. $75.50 more

OpenStudy (anonymous):

I would like to know if this is correct at all , Math is not my best subject.

OpenStudy (anonymous):

I will give you a medal <3

OpenStudy (anonymous):

Someone please help ;-; <3

OpenStudy (geneticrockhopper247):

It's been a long time since I've done this, so I might be wrong, but I will do the best that I can. You are charged interest on a mortgage at a bank and they pay you to keep your money there and interest is a percentage rate, so I'm going to guess that interest is all of the above on question 1. C on the second one because if you loan someone $100 and charge 5% interest, you get back $105. Do these two make sense?

OpenStudy (geneticrockhopper247):

On 3, $1,000 * 3.25% (aka .0325)= $32.50, so, yes, you are correct. On 4, $1,000*.0375= $37.50. 1,000 *.045= $45. 45-37.50=7.50, so the answer there would be C.

OpenStudy (geneticrockhopper247):

If there is anything here that doesn't make sense, please let me know.

OpenStudy (amistre64):

try 2 again

OpenStudy (amistre64):

5% interest on 100 is not 105

OpenStudy (amistre64):

100 + 5% interest = 105 how do we calculate the interest ....

OpenStudy (geneticrockhopper247):

Oh, I think I see where you're going.

OpenStudy (geneticrockhopper247):

my bad

OpenStudy (geneticrockhopper247):

Calculating interest, you would multiply the percentage against the total...I was thinking about using interest...

OpenStudy (geneticrockhopper247):

Actually, (pardon me and my schizoid answers), you would take the percentage and turn it into a decimal and multiply the amount by the decimal.